OpenStudy (anonymous):

**Correct Me :D [Work is Shown] Jennie has a jar of 54 coins consisting of dimes and quarters. If she has a total of $11.10, how many dimes does she have? 10d + 25(54) + 25(-d) =1110 10d + 1350 + 25(-d) =1110 10d + 1350 - 25d = 1110 -15d + 1350 = 1110 Subtract 1350. -15d = -240 Divided by -15 d = 16

OpenStudy (nurali):

d+q=54; 10d+25q=1110 d=16 q=38

OpenStudy (anonymous):

Looks great to me, but make sure you establish before you start your work what you are letting your variables represent. It's not always clear. I would mention that you let d=number of dimes and 54-d=number of quarters, for example.

OpenStudy (whpalmer4):

so you think she has 16 dimes and 54-16 = 38 quarters. let's see if that satisfies the problem 16*10 = 160 38*25 = 950 950+160 = 1110 your answer is correct.